### v1

@@ Abstract
Although Large Language Models (LLMs) show promising solutions to automated code generation, they often produce insecure code that threatens software security. Current approaches (e.g., SafeCoder) to improve secure code generation suffer from limited and imbalanced datasets, reducing their effectiveness and generalizability. In this work, we present Secure-Instruct, a novel framework that automatically synthesizes high‑quality vulnerable and secure code examples, generates fine‑tuning instructions, and instruction‑tunes LLMs to align task description and secure code generation abilities.
We evaluate Secure-Instruct on four representative LLMs using two benchmarks: our own CWEBench and the existing CWEval. CWEBench comprises 93 scenarios on 44 CWEs, all without overlap with Secure-Instruct’s synthetic instruction‑tuning dataset, while CWEval covers 31 CWEs with 119 manually verified security‑critical tasks. We find that Secure-Instruct improves not only the security but also the functional correctness of the generated code. On CWEBench, Secure-Instruct substantially improves secure code generation, giving a 14.3% average increase in secure ratio over the pretrained models and outperforms SafeCoder by 7.6%. On CWEval, Secure-Instruct achieves a 14% increase for CodeLlama-7B and 5.8% for Mistral-7B in Func-Sec@1 over pretrained models, and surpasses SafeCoder by 15.8% and 6.8% respectively.
@@ 1 Introduction
Large Language Models (LLMs), trained on billions of parameters across extensive corpora, have significantly advanced various software engineering tasks, such as code generation (Wei et al., 2024; Luo et al., 2023; Lin et al., 2024) and program repair (Jin et al., 2023; Joshi et al., 2023). Prominent models such as ChatGPT (OpenAI, 2024), Llama (Touvron et al., 2023), and GitHub Copilot (GitHub, 2024) have recently gained widespread popularity for their effectiveness in these domains.
@@
Despite the success of LLMs in generating functionally correct code, researchers have identified quality concerns in LLM-generated code. One of the most critical issues is the presence of security vulnerabilities. Pearce et al. (Pearce et al., 2022) revealed that approximately 40% of code snippets generated by GitHub Copilot contain vulnerabilities. Recent studies (Liu et al., 2024b; Sandoval et al., 2023) also found similar security issues in code generated by other LLMs. These findings suggest that, despite advances, vulnerabilities continue to pose a significant challenge in LLM-based code generation.
Several efforts have been made to enhance the secure code generation capabilities of LLMs. Prompt-based approaches (Hajipour et al., 2023; Wang et al., 2023; He and Vechev, 2023) design and experiment with different security-aware prompts to guide LLMs in generating secure code. However, prompt-based approaches struggle to scale and generalize across a wide variety of CWEs, leading to inconsistent improvements and generalization issues. Alternatively, fine-tuning LLMs (He et al., 2024; Li et al., 2024a, b) has been explored to improve LLMs’ secure code generation capabilities. Among all, SafeCoder (He et al., 2024) is the state-of-the-art (SOTA), which leverages real vulnerability fixes for fine-tuning LLMs and outperforms all prompt-based approaches.
@@
Yet, techniques like SafeCoder face two major limitations. First, exploiting real-world vulnerabilities and their fixes is limited by scale due to their rarity. For instance, SafeCoder managed to collect only 465 fixing examples covering 23 different CWEs with 6 programming languages from 145 million commits. This not only highlights the inefficiency of the approach but also its limited scalability, as even an extensive dataset contains examples for only a small subset of CWEs. Second, the methodology relies on a non-trivial manual inspection in the post-processing phase to balance and clean the dataset. Specifically, it down-samples over-represented CWE cases and manually removes examples that were incorrectly collected by the automated pipeline. Consequently, only 465 high-quality examples were retained from the 1,211 initially collected instances.
@@
To address the abovementioned challenges, we propose Secure-Instruct, a novel framework designed to synthesize instruction tuning datasets using LLMs from descriptive documentation of CWEs. Specifically, we utilize the CWE descriptions and code examples provided by the MITRE MITRE (2025) database as a foundational resource to prompt LLMs to generate diverse, high-quality, and secure code examples that LLMs can easily understand for later fine-tuning. Secure-Instruct integrates static security detection tools (SonarQube and CodeQL) to automatically ensure the high quality of LLM-synthesized datasets. In addition, inspired by recent works in instruction tuning (Longpre et al., 2023), our proposed Secure-Instruct leverages instruction tuning to better align LLMs with secure code generation. Secure-Instruct (1) bridges the gap between CWE documentation and practical secure code generation, (2) improves the performance of secure code generation in LLMs, (3) requires no manual efforts in curating high-quality fine-tuning datasets, and (4) can be easily extensible to work on new CWEs as it does not rely on open-source repositories for curating high-quality fine-tuning dataset for new CWEs.
Secure-Instruct experiments two synthetic data generation schemes: (1) vulnerable code driven and (2) secure-code driven. For (1), given a CWE description document, Secure-Instruct prompts an LLM to generate vulnerable code first, then continuously fixes the vulnerable code to achieve secure code, resulting in code pairs of (vul, secure). For (2), Secure-Instruct prompts an LLM to generate secure code practices directly (secure). Then, Secure-Instruct leverages LLMs to generate an instruction for each code snippet. Last, Secure-Instruct employs two instruction tuning strategies: one is the standard one Wei et al. (2024), and the other one is the secure fine-tuning approach proposed by He et al. (He et al., 2024). This approach uses a fine-tuning technique designed to learn secure coding patterns and assign penalties to vulnerable code snippets.
@@
We evaluate Secure-Instruct on four representative LLMs (CodeLlama-7B, CodeGen2-7B, Mistral-7B, and StarCoder-1B), which are commonly used in prior works on LLM secure code generation (He and Vechev, 2023; He et al., 2024) for fair comparison and are suitable under limited computing resources. Our evaluation includes two benchmarks, CWEBench (94 security scenarios covering 44 CWEs), and CWEval Peng et al. (2025) (119 manually verified security-critical tasks covering 31 CWEs). Our evaluation shows that the LLM fine-tuned by Secure-Instruct generates more secure code (an average improvement of 14.3%) compared to the pre-trained LLMs. On CWEBench, Secure-Instruct outperforms the state-of-the-art secure code generation approach, SafeCoder (He et al., 2024) by a large margin, +7.6% in CWEBench. On CWEval, Secure-Instruct achieves a better performance with +14.3% of Func-Sec@1 (both secure and functionally correct). Compared with the SOTA approach (SafeCoder), Secure-Instruct achieved +11.3% improvement in Func-Sec@1.
Furthermore, we performed ablation studies to understand the impact of fine-tuning datasets and approaches. Results show that the secure performance between Secure-Instructsecure and Secure-Instruct(vul, secure) is similar for the four models. Finally, we find that Secure-Instruct not only preserves but also improves LLMs’ capabilities in generating functionally correct code. The Pass@1 accuracy improved by +10.1% on CodeLlama-7B, +2.9% on Mistral-7B, and +1% on CodeGen2-7B based on the HumanEval benchmark (Chen et al., 2021). These results highlight the effectiveness of Secure-Instruct in advancing both secure and functional code generation.
The main contributions of this work include:
  • We propose Secure-Instruct, a fully automated framework that synthesizes high-quality datasets from CWE documentation to instruction tune LLMs for secure code generation. Secure-Instruct is generalizable and scalable to a diverse set of CWEs and is completely free of manual effort.
  • We extend existing secure benchmarks to a total of 93 security-sensitive prompts covering 44 CWEs to provide a comprehensive evaluation on secure code generation.
@@
  • We conduct a comprehensive evaluation for Secure-Instruct on four LLMs, and we take a comparative analysis of secure performance with the SOTA approach. Our approach not only significantly improves the secure code generation but also improves the functionality correctness evaluated on CWEval, HumanEval, and MBPP benchmarks.
  • We share the synthesized fine-tuning dataset, the replication code, and fine-tuned models’ weights at https://zenodo.org/records/15019488
@@ 2 Background and Related Work
Security Concerns of LLM-Generated Code. Ensuring program security remains a challenge in both industry and academia. While LLMs have demonstrated significant potential in automating code generation, they are often trained on unsanitized datasets from open-source repositories, which can result in the generation of insecure code. Figure 1 shows an example of vulnerable code generated by CodeLlama-7B. This vulnerable code (lines 10 and 11) demonstrates a CWE-78 (‘OS Command Injection’). ‘username’ comes directly from ‘argv[1]’, which is not sanitized or validated. An attacker could inject arbitrary commands by passing a specially crafted username as input.
Several studies (Sandoval et al., 2023; Siddiq and Santos, 2023; Gong et al., 2024; Tihanyi et al., 2025; Fu et al., 2023; Mohsin et al., 2024) have investigated the security vulnerabilities present in LLM-generated code. Pearce et al. (Pearce et al., 2022) evaluated GitHub Copilot, which is built on the OpenAI Codex model, and reported that approximately 40% of the generated code contained security vulnerabilities. These vulnerabilities are systematically classified under the Common Weakness Enumeration (CWE) (CWE-MIRTE, 2022), a widely adopted category system for security vulnerabilities. Khoury et al. (Khoury et al., 2023) examined ChatGPT in security-critical coding scenarios, discovering that the model produced insecure code in 16 cases, with only 7 cases corrected after iterative interaction with the model. Similarly, Perry et al. (Perry et al., 2023) observed that developers relying on AI-assisted tools are more likely to introduce security vulnerabilities.
@@
Improving LLMs for Better Code Generation. Acquiring an effective and high-quality fine-tuning dataset from real-world sources is often expensive and labor-intensive. To address this challenge, researchers have explored methods for synthesizing datasets using large language models (LLMs). Recent advancements have introduced synthetic data generation techniques during fine-tuning to address data acquisition challenges and enhance LLM performance on specific tasks (Babbar and Schölkopf, 2019; Long et al., 2024; Gilardi et al., 2023; Chang et al., 2024; Liu et al., 2024a; Abay et al., 2019).
Self-Instruct Wang et al. (2022) introduced an automated approach for instruction tuning dataset generation using LLM. Subsequent works extend this strategy to a broader range of models and tasks Taori et al. (2023); Xu et al. (2024), including code generation Chaudhary (2023); Luo et al. (2023); Wei et al. (2024). Evol-Instruct (Xu et al., 2024; Luo et al., 2023) uses ChatGPT to generate diverse instruction-following examples using carefully crafted prompts. Wei et al. Wei et al. (2024) proposed OSS-Instruct, a method that uses ChatGPT to generate diverse coding problems and solutions from seed code snippets from open-source code, to create a high-quality fine-tuning dataset to improve code generation accuracy.
Improving the Security of LLM Generated Code.
To address security problems in LLM-generated code, Li et al. Li et al. (2024b) proposed an approach to fine-tune one LLM by using the vulnerability-fixing commits. Hajipour et al. (Hajipour et al., 2024) introduced a technique to synthesize secure code examples guided by CodeQL (CodeQL, 2024) oracle. Another work (Li et al., 2024a) explored the effect of parameter-efficient fine-tuning techniques in secure code generation. He et al. (He and Vechev, 2023) introduced SVEN, an advanced prompting method, i.e., prefix-tuning, that uses property-specific continuous vectors (prefixes) to guide LLMs to generate secure code and reduce the likelihood of generating vulnerable code. The training of SVEN optimizes these continuous vectors using a carefully curated, high-quality dataset.
The most recent work, SafeCoder (He et al., 2024), proposes a security-focused instruction tuning approach that learns from vulnerable–fix code pairs mined from 145 million commits. Through semi-automated filtering, the authors curate a dataset of 465 high-quality code examples. Their findings show that the LLM fine-tuned on this data achieves state-of-the-art performance in secure code generation.
Despite improved results, SafeCoder has several limitations. First, the quality of code from open-source repositories is inconsistent, often involving complex dependencies with external libraries. As a result, many code snippets are unsuitable for fine-tuning, requiring significant manual effort to filter and curate appropriate samples. Second, it is challenging to obtain a sufficiently large fine-tuning dataset for each category of security issues (e.g., CWE). For example, the work (He et al., 2024) identified 71 code snippets for CWE-079 but only one for CWE-732, resulting in an imbalance that hinders the model’s ability to achieve robust security performance. Besides, this approach is expensive. They only collected 456 code snippets covering 23 CWEs from 145 million commits, making it difficult to expand to other CWEs. To address these limitations, we propose Secure-Instruct to automatically construct a more extensive and diverse dataset to further enhance the security of LLM-generated code.
@@ 3 The Methodology of Secure-Instruct
We introduce Secure-Instruct, a framework that leverages LLMs (e.g., ChatGPT and Claude) to synthesize fine-tuning data for instruction-tuning code language models (CLMs) to enhance secure code generation. Figure 2 shows an overview of Secure-Instruct. It adopts two data-generation schemes, secure-code driven (Secure-Instructsecure) and vulnerable-code driven (Secure-Instruct(vul, secure)), to generate high-quality secure training data with no human intervention (Section 3.2). It then uses LLMs to generate natural-language instructions that describe the functionality of each code example in the dataset. Finally, Secure-Instruct applies two instruction tuning strategies to produce security-enhanced LLMs (Section 3.3): standard instruction tuning (Longpre et al., 2023) and security-specialized instruction tuning (He et al., 2024).
@@ 3.1 CWE Seeds Collection
First, we collect CWE documents from the MITRE. The MITRE is an official website that serves as an authoritative resource for understanding and analyzing Common Weakness Enumerations (CWEs). Each CWE entry on the website includes a comprehensive description, highlighting the nature of the vulnerability, its potential impact, and common scenarios in which it arises. Additionally, the entries feature illustrative code examples that contrast vulnerable and secure implementations, offering both context and practical insights.
Figure 3 presents an example of CWE-78 (OS Command Injection) from the MITRE website. This example includes a detailed overview of the weakness, an explanation of its root causes, and multiple code snippets demonstrating how this vulnerability can manifest in real-world software. These examples not only facilitate a deeper understanding of the security flaws but also serve as valuable references for LLMs to generate secure code examples. In particular, Secure-Instruct leverages the information from the MITRE website for each CWE, namely 1) the category of CWE, 2) the overall description, and 3) code snippet examples with their explanations. In total, we collected 44 CWEs from MITRE.
@@ 3.2 Synthetic Data Generation By Prompting LLMs
Given seed CWE documents, Secure-Instruct employs two data generation strategies to prompt LLM for generating fine-tuning datasets, namely vulnerable-code driven (Secure-Instruct(vul, secure)) and secure-code driven (Secure-Instructsecure).
@@
Secure-Instruct(vul, secure). As Figure 2 illustrates, vulnerable-code driven data generation includes four steps: (1) prompting an LLM (e.g., GPT 4o) to generate vulnerable code given one CWE information; (2) verifying that the LLM-generated code indeed contains the specified CWE; (3) prompting LLMs to fix the vulnerabilities and produce secure code; and (4) confirming the secure code is vulnerability-free. Vulnerable-code driven generation ensures the success of generating both vulnerable and secure code for specific CWEs. We employ two powerful LLMs for code generation, i.e., ChatGPT (gpt-4o-2024-08-06) and Claude (claude-3-5-sonnet-20241022), to generate intentionally vulnerable code snippets. Below, we show the prompt templates for generating and fixing vulnerable code.
Prompt Template for Vulnerable Code Generation of Secure-Instruct(vul, secure). Task: Please generate an easily understandable vulnerable code snippet based on the following description of CWE. (Please consider the diversity of the vulnerable code examples for each generation.) CWE: {<CWE_OVERALL_DESCRIPTION>} A Vulnerable Code Example: This is {<LANGUAGE>} language example.
{<CODE>}
{<GENERATED_VULNERABLE_CODE>}
Can you FIX the code for {<LANGUAGE>} ? Please make sure the code is secure to the CWE mentioned above and runnable.
For each pair of “CWE-ID” and “programming language”, Secure-Instruct(vul, secure) takes both the CWE description and a randomly selected code example from the seed corpus and prompts each LLM to generate 10 vulnerable code snippets. In total, Secure-Instruct(vul, secure) collected 1,560 code snippets spanning 78 CWE-language pairs. Then, Secure-Instruct(vul, secure) utilizes static security analyzers to verify the vulnerabilities present in the generated code. Finally, we instruct the LLMs to fix the vulnerable code and verify that the vulnerabilities are gone using static security analyzers.
Since static analysis tools can produce both false positives and false negatives, we use two static security analyzers, CodeQL (CodeQL, 2024) and SonarQube (SonarQube, 2025), to complement each other in verification. A code snippet is considered vulnerable only if both tools produce identical detection results. However, while CodeQL supports all 78 CWE-language pairs, SonarQube only supports a subset of 41 pairs (i.e., 6 CWEs for C, 9 for Java, 12 for Python, 9 for JavaScript, 0 for Ruby, and 5 for Go). For the cases not supported by SonarQube, Secure-Instruct solely relies on CodeQL.
@@
After verification of the 1,560 snippets by static analyzers, 475 code snippets were confirmed to be truly vulnerable. These verified snippets then served as input seeds for the LLMs to generate five fixes for each snippet. Finally, static analyzers are used again to verify whether the generated fixes were valid. In the end, Secure-Instruct(vul, secure) collected 856 pairs of vulnerable code and their corresponding fixed code, which cover 25 CWEs in 6 programming languages, i.e., 39 CWE-language pairs.
@@
Secure-Instructsecure. Secure-Instruct(vul, secure)’s double verification process results in many LLM-generated code snippets being filtered out, e.g., LLMs may fail to generate vulnerable code or fix vulnerable code in many cases, leading to a limited set of secure code snippets. To source an even larger set of secure code snippets, we propose an alternative approach, Secure-Instructsecure. Instead of generating vulnerable code and then repairing it, we directly instruct LLMs to produce secure code using CWE documents as prompts. For each CWE-language pair, we generate 100 secure code snippets that pass verification by both CodeQL and SonarQube, ensuring they do not contain vulnerabilities.

### v2

@@ Abstract
Although Large Language Models (LLMs) show promising solutions to automated code generation, they often produce insecure code that threatens software security. Current approaches (e.g., SafeCoder) to improve secure code generation are limited by small, imbalanced instruction-tuning datasets. In this work, we present Secure-Instruct, a novel pipeline that automatically synthesizes high‑quality vulnerable and secure code examples and instruction-tunes LLMs to align task description and secure code generation abilities.
We evaluate Secure-Instruct on four representative LLMs using two security-related benchmarks: our own CWEBench and the existing CWEval. CWEBench comprises 93 scenarios on 44 CWEs, all without overlap with Secure-Instruct’s synthetic instruction‑tuning dataset, while CWEval covers 31 CWEs with 119 manually verified security‑critical tasks. We find that Secure-Instruct improves both security and functional correctness in code generation. On CWEBench, Secure-Instruct substantially improves secure code generation, giving a 28.5% increase on average in secure ratio over the pre-trained models and outperforms SafeCoder by 12.6%. On CWEval, Secure-Instruct achieves an increase of 157.3% for CodeLlama-7B and 46.4% for Mistral-7B in Func-Sec@1 over pretrained models, and significantly outperforms SafeCoder.
@@ 1 Introduction
Large Language Models (LLMs), trained on billions of parameters across extensive corpora, have significantly advanced various software engineering tasks, such as code generation (wei2024magicoder; luo2023wizardcoder; soen101) and program repair (jin2023inferfix; joshi2023repair). Prominent models such as ChatGPT (openai2024chatgpt), Llama (touvron2023llama), and GitHub Copilot (github2024copilot) have recently gained widespread popularity for their effectiveness in these domains.
@@
Despite their success, LLMs often generate code with security vulnerabilities. Pearce et al. (pearce2022asleep) revealed that approximately 40% of code snippets generated by GitHub Copilot contain vulnerabilities, and similar issues have been reported in other models (liu2024no; sandoval2023lost). These findings suggest that, despite advances, vulnerabilities continue to pose a significant challenge in LLM-based code generation.
@@
Several efforts (hajipour2023codelmsec; wang2023enhancing; he2023large; he2024instruction; li2024exploratory; li2024fine) have been made to enhance the secure code generation capabilities of LLMs. Among all, SafeCoder (he2024instruction) is the state-of-the-art (SOTA), which leverages real vulnerability fixes for instruction-tuning LLMs and outperforms other approaches. Yet, techniques like SafeCoder face two major limitations. First, exploiting real-world vulnerabilities and their fixes is limited by scale due to their rarity. For instance, SafeCoder managed to collect only 465 fixing examples covering 23 different Common Weakness Enumerations (CWEs) with 6 programming languages from 145 million commits. This not only highlights the inefficiency of the approach but also its limited scalability, as even an extensive dataset contains examples for only a small subset of CWEs. Second, the methodology relies on a non-trivial manual inspection in the post-processing phase to balance and clean the dataset. Specifically, it down-samples over-represented CWE cases and manually removes examples that were incorrectly collected by the automated pipeline. Consequently, only 465 high-quality examples were retained from the 1,211 initially collected instances.
@@
To address the abovementioned challenges, we propose Secure-Instruct, a novel and automated pipeline designed to synthesize instruction tuning datasets from descriptive documentation of CWEs. Specifically, we utilize the CWE descriptions and code examples provided by the MITRE (mitre-cwe) database as a foundational resource to prompt LLMs to generate diverse, high-quality, and secure code examples that LLMs can easily understand for later instruction-tuning. Secure-Instruct integrates static security detection tools (SonarQube and CodeQL) to automatically ensure the high quality of LLM-synthesized datasets. In addition, inspired by recent works in instruction tuning (longpre2023flan), our proposed Secure-Instruct leverages instruction tuning to better align LLMs with secure code generation. Secure-Instruct (1) bridges the gap between CWE documentation and practical secure code generation, (2) improves the performance of secure code generation in LLMs, (3) requires no manual efforts in curating high-quality instruction-tuning datasets, and (4) can be easily extensible to work on new CWEs as it does not rely on open-source repositories for curating high-quality instruction-tuning dataset for new CWEs.
@@
Secure-Instruct experiments with two synthetic data generation schemes: (1) vulnerable-code driven and (2) secure-code driven. For (1), given a CWE description document, Secure-Instruct prompts an LLM to generate vulnerable code first, then continuously fixes the vulnerable code to achieve secure code, resulting in code pairs of (vul, secure). For (2), Secure-Instruct prompts an LLM to generate secure code practices directly (secure). Then, Secure-Instruct leverages LLMs to generate an instruction for each code snippet. Last, Secure-Instruct employs two instruction tuning strategies: one is the standard one (wei2024magicoder), and the other is the secure instruction-tuning approach proposed by He et al. (he2024instruction). This approach uses an instruction-tuning technique designed to learn secure coding patterns and assign penalties to vulnerable code snippets. We evaluate Secure-Instruct on four representative LLMs (CodeLlama-7B, CodeGen2-7B, Mistral-7B, and StarCoder-1B), which are commonly used in prior works on LLM secure code generation (he2023large; he2024instruction) for fair comparison and are suitable under limited computing resources. Our evaluation includes two benchmarks, CWEBench (94 security scenarios covering 44 CWEs), and CWEval (peng2025cweval) (119 manually verified security-critical tasks covering 31 CWEs). On CWEBench, Secure-Instruct outperforms the state-of-the-art secure code generation approach, SafeCoder (he2024instruction) by a large margin, with an average improvement of 12.6%. On CWEval, Secure-Instruct achieves better performance with an average 56.6% improvement of Func-Sec@1 (both secure and functionally correct) compared to pre-trained models, and it outperforms SafeCoder with an average improvement of 140.8%. Finally, it also improves functional correctness with +39.6% (CodeLlama-7B), +12.6% (Mistral-7B), and +6.5% (CodeGen2-7B) on HumanEval (chen2021codex). These results highlight the effectiveness of Secure-Instruct in advancing both secure and functional code generation.
@@
Contributions. 1) We propose Secure-Instruct, a fully automated framework that synthesizes high-quality datasets from CWE documentation to instruction tune LLMs for secure code generation. Secure-Instruct is generalizable and scalable to a diverse set of CWEs and is completely free of manual effort. 2) We extend existing secure benchmarks to a total of 93 security-sensitive prompts covering 44 CWEs to provide a comprehensive evaluation on secure code generation. 3) We conduct a comprehensive evaluation for Secure-Instruct on four LLMs, and we take a comparative analysis of secure performance with the SOTA approach. Our approach not only significantly improves the secure code generation but also improves the functionality correctness evaluated on CWEval, HumanEval, and MBPP benchmarks. 4) We share the synthesized instruction-tuning dataset, the replication code, and models’ weights at https://zenodo.org/records/16715355.
@@ 2 Background and Related Work
Security Concerns of LLM-Generated Code. While LLMs show promise in automating code generation, their training on unsanitized open-source data often leads to insecure code. Figure 1 shows an example of vulnerable code generated by CodeLlama-7B. This vulnerable code (line 8) demonstrates a CWE-78 (‘OS Command Injection’). ‘username’ comes directly from ‘argv[1]’, which is not sanitized or validated. An attacker could inject arbitrary commands by passing a specially crafted username as input.
Several studies (sandoval2023lost; siddiq2023generate; gong2024well; tihanyi2025secure; fu2023security; mohsin2024can) have investigated the security vulnerabilities present in LLM-generated code. Pearce et al. (pearce2022asleep) evaluated GitHub Copilot, which is built on the OpenAI Codex model, and reported that approximately 40% of the generated code contained security vulnerabilities. These vulnerabilities are systematically classified under the Common Weakness Enumeration (CWE) (CWE), a widely adopted category system for security vulnerabilities. Khoury et al. (khoury2023secure) examined ChatGPT in security-critical coding scenarios, discovering that the model produced insecure code in 16 cases, with only 7 cases corrected after iterative interaction with the model. Similarly, Perry et al. (perry2023users) observed that developers relying on AI-assisted tools are more likely to introduce security vulnerabilities.
@@
Improving the Security of LLM Generated Code. To overcome these limitations, Li et al. (li2024fine) proposed an approach to fine-tune one LLM by using the vulnerability-fixing commits. Hajipour et al. (hajipour2024hexacoder) introduced a technique to synthesize secure code examples guided by CodeQL (codeql) oracle. Another work (li2024exploratory) explored the effect of parameter-efficient fine-tuning techniques in secure code generation. He et al. (he2023large) introduced SVEN, an advanced prompting method, i.e., prefix-tuning, that uses property-specific continuous vectors (prefixes) to guide LLMs to generate secure code and reduce the likelihood of generating vulnerable code. The training of SVEN optimizes these continuous vectors using a carefully curated, high-quality dataset.
Most recently, SafeCoder (he2024instruction) applies instruction-tuning on a dataset of 465 vulnerable-fix pairs, achieving state-of-the-art results in secure code generation. Despite improved results, SafeCoder has several limitations. First, the quality of code from open-source repositories is inconsistent, often involving complex dependencies with external libraries. As a result, many code snippets are unsuitable for fine-tuning, requiring significant manual effort to filter and curate appropriate samples. Second, it is challenging to obtain a sufficiently large fine-tuning dataset for each category of security issues (e.g., CWE). For example, the work (he2024instruction) identified 71 code snippets for CWE-079 but only one for CWE-732, resulting in an imbalance that hinders the model’s ability to achieve robust security performance. Besides, this approach is expensive. They only collected 456 code snippets covering 23 CWEs from 145 million commits, making it difficult to expand to other CWEs. To address these limitations, we propose Secure-Instruct to automatically construct a more extensive and diverse dataset to further enhance the security of LLM-generated code.
@@ 3 The Methodology of Secure-Instruct
We introduce Secure-Instruct that leverages LLMs (e.g., ChatGPT and Claude) to synthesize data for instruction-tuning code language models to enhance secure code generation. Figure 2 shows an overview of Secure-Instruct. It adopts two data-generation schemes, secure-code driven (Secure-Instructsecure) and vulnerable-code driven (Secure-Instruct(vul, secure)), to generate high-quality secure training data with no human intervention. It then uses LLMs to generate natural-language instructions that describe the functionality of each code example in the dataset. Finally, Secure-Instruct applies two instruction-tuning strategies to produce security-enhanced LLMs (Section 3.3): standard instruction tuning (longpre2023flan) and security-specialized instruction tuning (he2024instruction).
@@ 3.1 CWE Seeds Collection
First, we collect Common Weakness Enumeration (CWE) documents from the MITRE. The MITRE is an official website that serves as an authoritative resource for understanding and analyzing CWEs. Each CWE entry on the website includes a comprehensive description, highlighting the nature of the vulnerability, its potential impact, and common scenarios in which it arises. Additionally, the entries feature illustrative code examples that contrast vulnerable and secure implementations, offering both context and practical insights. As shown in Figure 3, CWE-78 (OS Command Injection) provides such insights. Secure-Instruct leverages the information from the MITRE website for each CWE, namely 1) the category of CWE, 2) the overall description, and 3) code snippet examples with their explanations. In total, we collected 44 CWEs from MITRE.
@@ 3.2 Synthetic Data Generation By Prompting LLMs
Given seed CWE documents, Secure-Instruct employs two data generation strategies to prompt LLM for generating instruction-tuning datasets, namely vulnerable-code driven (Secure-Instruct(vul, secure)) and secure-code driven (Secure-Instructsecure).
@@
Secure-Instruct(vul, secure). As Figure 2 illustrates, the vulnerable-code driven data generation includes four steps: (1) prompting an LLM (e.g., GPT 4o) to generate vulnerable code given one CWE information; (2) verifying that the LLM-generated code indeed contains the specified CWE; (3) prompting LLMs to fix the vulnerabilities and produce secure code; and (4) confirming the secure code is vulnerability-free. Vulnerable-code driven generation ensures the success of generating both vulnerable and secure code for specific CWEs. We employ two powerful LLMs for code generation, i.e., ChatGPT (gpt-4o-2024-08-06) and Claude (claude-3-5-sonnet-20241022), to generate intentionally vulnerable code snippets. The templates for generating and fixing vulnerable code are included in our replication package. For each pair of “CWE” and “programming language”, Secure-Instruct(vul, secure) takes both the CWE description and a randomly selected code example from the seed corpus and prompts each LLM to generate 10 vulnerable code snippets. In total, Secure-Instruct(vul, secure) collected 1,560 code snippets spanning 78 CWE-language pairs. Then, Secure-Instruct(vul, secure) utilizes static security analyzers to verify the vulnerabilities present in the generated code. Finally, we instruct the LLMs to fix the vulnerable code and verify that the vulnerabilities are gone using static security analyzers. Since static analysis tools can produce both false positives and false negatives, we use two static security analyzers, CodeQL (codeql) and SonarQube (sonarqube), to complement each other in verification. A code snippet is considered vulnerable only if both tools produce identical detection results.
@@
After verification of the 1,560 snippets by static analyzers, 475 were confirmed to be truly vulnerable. These verified snippets then served as input seeds for the LLMs to generate five fixes for each snippet. Finally, static analyzers are used again to verify whether the generated fixes were valid. After generating the dataset through our pipeline and verifying it using static analyzers, we performed a de-duplication step to eliminate potential overlaps with evaluation benchmarks. This ensures that our evaluation remains fair and unbiased. The details of this de-duplication process are described in Section 4. In the end, Secure-Instruct(vul, secure) produced 856 verified (vulnerable, fixed) pairs across 25 CWEs and 6 languages, covering 39 CWE-language combinations.
@@
Secure-Instructsecure. Secure-Instruct(vul, secure)’s double verification process results in many LLM-generated code snippets being filtered out, e.g., LLMs may fail to generate vulnerable code or fix vulnerable code in many cases, leading to a limited set of secure code snippets. To source an even larger set of secure code snippets, we propose an alternative approach, Secure-Instructsecure. Instead of generating vulnerable code and then repairing it, we directly instruct LLMs to produce secure code using CWE documents as prompts. For each CWE-language pair, we generate 100 secure code snippets that pass verification by both CodeQL and SonarQube, ensuring they do not contain vulnerabilities. Similarly, we performed a de-duplication step to make sure that all generated code snippets are not duplicated with evaluation benchmarks.